OpenStudy (anonymous):

Simplify. Express the product as a rational exponent. x to the 1 third power times x to the 1 third power

OpenStudy (kira_yamato):

x^(2/3)

OpenStudy (turingtest):

[x^(1/3)][x^(1/3)]=x^(1/3+1/3)=x^(2/3)

OpenStudy (anonymous):

\[x^{1/3} \times x^{1/3} =x^{1/3+1/3}=x^{2/3}\] It tooks me forever to type it right way